If a population with size 3 has a standard deviation of 2.4, what is thestandard deviation of the sampling distribution of its means? The sampling distribution has a sample size of 2 and all possible samples are drawn without replacements. 


1
Expert's answer
2021-05-11T09:07:02-0400

We have given,


"n = 3"


"\sigma = 2.4"


"\sigma_{\bar x} = \dfrac{\sigma}{\sqrt{n}}"


"\sigma_{\bar x} = \dfrac{2.4}{\sqrt{3}}"


"\sigma_{\bar x} = 1.38"